# testphase=new-or-1

### v1

@@ Abstract
Reward models (RMs) are central to aligning large language models (LLMs) with human values but have received less attention than pre-trained and post-trained LLMs themselves. Because RMs are initialized from LLMs, they inherit representations that shape their behavior, but the nature and extent of this influence remain understudied. In a comprehensive study of 10 leading open-weight RMs using validated psycholinguistic corpora, we show that RMs exhibit significant differences along multiple dimensions of human value as a function of their base model. Using the “Big Two” psychological axes, we show a robust preference of Llama RMs for “agency” and a corresponding robust preference of Gemma RMs for “communion.” This phenomenon holds even when the preference data and finetuning process are identical, and we trace it back to the logits of the respective instruction-tuned and pre-trained models. These log-probability differences themselves can be formulated as an implicit RM; we derive usable implicit reward scores and show that they exhibit the very same agency/communion difference. We run experiments training RMs with ablations for preference data source and quantity, which demonstrate that this effect is not only repeatable but surprisingly durable. Despite RMs being designed to represent human preferences, our evidence shows that their outputs are influenced by the pretrained LLMs on which they are based. This work underscores the importance of safety and alignment efforts at the pretraining stage, and makes clear that open-source developers’ choice of base model is as much a consideration of values as of performance.
@@ 1 Introduction
Reward models (RMs) play a key role in aligning large language models (LLMs) with human preferences and values. Reward modeling can be “explicit,” relying on a reinforcement learning–based approach for learning from human feedback (RLHF; Christiano et al. 2017), or “implicit,” directly increasing the probability of human-preferred data through a cross-entropy objective (Rafailov et al., 2023). Despite their central importance in AI safety, RMs have received relatively less attention than both pre-trained and post-trained LLMs. This has recently started to change with the increased availability of human preference data (Bai et al., 2022; Liu et al., 2024; Jiang et al., 2023), of open-weight RMs, and of public RM benchmarks (Lambert et al., 2024; Malik et al., 2025). Recent work on RM interpretability has focused on how RMs may be used to intentionally bias post-trained models towards specific preferences – e.g., model personalization (Luo et al., 2025; Wang et al., 2024; Sorensen et al., 2024) – or on how RMs may unintentionally introduce bias in post-trained LLMs (Siththaranjan et al., 2023; Bharadwaj et al., 2025; Kumar et al., 2025). However, RMs are typically initialized from LLMs before being finetuned for preference modeling, and no work to date has looked at how RMs themselves can be biased by the LLMs from which they are built. This is a particularly worrying knowledge gap in light of recent research highlighting the importance of pretraining choices in model misalignment (Maini et al., 2025; O’Brien et al., 2025; Chen et al., 2025b). Given RMs’ key role in alignment pipelines, it is crucial to understand their vulnerability to potential sources of value bias from pretraining.
@@
In this paper, we systematically investigate whether RMs inherit value biases from pretraining. We use the “exhaustive token search” method introduced by Christian et al. (2025), in which RM reward scores are obtained across the entire token vocabulary to reveal the highest- and lowest-scoring responses to user prompts, and we combine this approach with tools from psycholinguistics (Pennebaker et al., 2003) to uncover and quantify value biases in RMs as a function of the base model on which they are developed. We analyze data from 10 leading RMs on RewardBench and find robust and replicable differences between Llama- and Gemma-based RMs across a variety of dimensions of human value (Section 2). As a case study, we focus on the Big Two psychological dimensions (Bakan, 1966; Abele and Wojciszke, 2018) that capture agency-oriented values (e.g., freedom, success, ability) and communion-oriented ones (e.g., love, family, friendship). We use a psychologically-validated corpus of words relating to agency and communion to demonstrate a robust relative preference by Llama-based RMs for agency, and by Gemma-based RMs for communion. Next, we trace the source of those biases to the base models themselves (Section 3) and explore differences between the Llama and Gemma base models, as implied by differences in their logprobs (relating to implicit reward models). Finally, we conduct systematic experiments training our own RMs on different base models with identical data and hyperparameters, using various sources and ablations of data, in order to chart how the observed bias evolves over the course of preference finetuning and the extent to which it can – or cannot – be “washed out” with sufficient finetuning data (Section 4).
@@
    We trace these differences back to differences in the log probabilities of the instruction-tuned models, and ultimately, in the pre-trained models on which the RMs are built.
@@ 2 RMs in the Wild Show Value Differences by Base Model
Figure 1: Value preferences (token ranks) from 10 leading RewardBench RMs based on Gemma and Llama for words related to different moral concepts. (a) Preferences for the Big-Two dimensions, for positively-framed prompts (top) and negatively-framed prompts (bottom). (b) Same as (a), for 5 MFD2 dimensions. Dots show mean ± s.e. of the median ranking of each single model, averaged over prompts; black markers indicate grand mean ± s.e.; violin plots visualize the density of the distribution. Stars: $p<.0001$ (Bonferroni-corrected permutation $t$-tests).
We evaluate the rank-ordered reward scores assigned by the same set of 10 leading Gemma- and Llama-based RMs from RewardBench (list in Appendix A) to words contained in the Big2 and MFD corpora as responses to a set of 54 value-laden prompt variations (details in Appendix E). The resulting dataset comprises 263 (Big Two) or 2,040 (MFD2) word rankings $\times$ 10 models $\times$ 54 prompts (27 of which were positively framed, e.g., “What, in one word, is the greatest thing ever?” and 27 of which were negatively framed, e.g., “the worst thing ever”). We quantify the effect of base model on the median rank assigned to words from each value category via a mixed-effects linear model, where we include fixed effects for prompt variation and interactions with value category, and group data by each individual RM (each individual data point in Fig. 1 represents a single RM; in Appendix C we visualize all prompt-model pairs).
@@
In positively framed prompts, Llama RMs rank agency-related words (including “success,” “skills,” “capability”) more highly than Gemma RMs, and Gemma-based RMs rank communion-related words (including “love,” “friends,” “relationships”) higher than Llama-based RMs. The opposite is true for negative prompts: Llama RMs prefer communion words (as answers to “the worst thing”) relative to Gemma, and Gemma RMs prefer agency words relative to Llama (3-way interaction between Big-Two category $\times$ base model $\times$ prompt valence, $p<.0001$, all follow-up permutation-based $t$-tests, $p<.0001$). These differences between base models constitute a medium effect size, Cohen’s $d$ of .40–.43.
@@
In positively framed prompts, Llama RMs rank authority- and fairness-related words better compared to Gemma, and Gemma RMs rank care-, loyalty- and sanctity-related words higher than Llama (permutation-based $t$-tests, all $p<.0001$). For the negatively framed prompts, the results are less clear cut. We find the (expected) opposite pattern for care (Llama >> Gemma, $p<.0001$), suggesting that Gemma prioritizes care-related words relative to Llama. However, for authority, loyalty and sanctity the pattern was the same for positive and negative prompts (all $p<.0001$); the fairness contrast did not reach our Bonferroni-corrected criterion alpha level of $p=0.00125$.
These results indicate that choice of base model significantly impacts rankings of words relating to different dimensions of value. We find consistent evidence (see Appendix D for reproduction of these results with existing data from Christian et al. (2025)’s exhaustive token search) that RMs based on Llama and Gemma exhibit biases toward agency and communion, respectively, and differ along a variety of other axes of value. We take the clear agency/communion finding as a case study to trace both the pre-trained origins of these biases in Section 3 as well as their evolution during reward modeling in Section 4.
3 Value Biases Begin in Pre-training
If the RMs analyzed in Section 2 inherited their biases from their base models, then we should expect to observe a similar bias in the instruction-tuned versions of Gemma and Llama on which those RMs are based – and likely also in the pre-trained Gemma and Llama models on which those are based. We investigated these Gemma and Llama LLMs using two different methods: looking directly at the models’ individual log probabilities, as well as computing a metric that is able to represent the difference between the two LLM policies as an implicit reward model itself. In both cases, we find precisely the phenomenon that we observed in the behavior of the downstream RMs, revealing that the effect reported in Section 2 is, indeed, rooted in the base models themselves.
@@ 3.1 Log Probabilities Mirror RM Agency/Communion Biases
Using the same set of prompts as in Section 2, we calculated the log probability assigned to each Big-Two noun by the instruction-tuned versions of Gemma 2 2B and Llama 3.2 3B. Fig. 2 shows the median rank of agency and communion words. Consistent with the pattern observed in the RMs, we find that in positively framed prompts, agency words are ranked higher by Llama, while communion words are ranked higher by Gemma. This pattern is reversed for the negatively-framed prompts. A three-way ANOVA revealed a significant interaction between Big-Two category, prompt valence, and model ($F(1,208)=88.8$, $p<0.0001$). We find the same interaction in the pre-trained versions of Gemma 2 2B and Llama 3.2 3B ($F(1,208)=42.3$). Welch’s $t$-tests for all relevant comparisons yielded FDR-corrected $p<0.01$. This analysis is carried out on the subset of 82 Big-Two nouns (lowercase) that are present in both Gemma and Llama tokenizer vocabularies.
@@
(b) Pre-trained
Figure 2: Log probabilities in both the instruction-tuned and pre-trained versions of the Gemma and Llama base models reveal the same agency/communion split observed in their respective RMs’ reward scores. Violin plots show the median rank of the Big-Two nouns according to the log probabilities assigned by the (a) instruction-tuned and (b) pre-trained versions of Gemma 2 2B and Llama 3.2 3B. Each dot corresponds to one of our positively (top) or negatively (bottom) valenced prompts. *** $p<0.001$, ** $p<0.01$, FDR-corrected. Boxes show median (white line) and interquartile ranges and whiskers extend to the ends of the distribution excluding outliers.
@@ 3.2 Implicit Reward Scores Mirror RM Agency/Communion Biases
where $r(x,y)$ is the reward for the concatenated sequence $[x,y]$. In practice, this is achieved solving a regularized RL problem to which $\pi_{\text{r}}$} is the solution:
$\pi_{\text{r}}(y|x)=\arg\max_{\pi}\mathbb{E}_{x\sim\pi}[r(x)]-\frac{1}{\beta}KL[\pi||\pi_{\text{base}}].$
@@
$r_{1\to 2}(x,y)=c(x)+\beta\cdot\log\frac{\pi_{\text{2}}(y|x)}{\pi_{\text{1}}(y|x)}$
Hence, for a given prompt $x$, the log difference $\log\,{\pi_{\text{2}}(y|x)}-\log\,{\pi_{\text{1}}(y|x)}$, can be interpreted as a relative implicit reward, on top of which an “exhaustive token search” methodology may be applied to reveal “optimal” and “pessimal” tokens.
@@
To address this problem, we considered several alternative measures designed to avoid spurious contributions from low-probability tokens. Letting $p(\cdot)\equiv\pi_{1}(\cdot\mid x)$, and $q(\cdot)\equiv\pi_{2}(\cdot\mid x)$, a particularly natural choice is to weight the log-probability difference by the probability of the token under the mixture:
@@
To evaluate the empirical usefulness of the MWLR score against other candidate scores, we create an “authoritarian” version of Gemma 2 IT 2B by boosting 10 words from the MFD “authority.virtue” list via supervised finetuning, and then inspect which candidate measures are best able to recover those words. The MWLR score outperforms all other measures tested in sensitivity to the induced value shifts (details in Appendix F).
@@
Equipped with a usable implicit-RM score, we use it to characterize the values that distinguish Gemma from Llama. What implicit RM, if given Gemma 2 2B as a base model to finetune, would produce Llama 3.2 3B? And what would be the “optimal and pessimal tokens” Christian et al. (2025) for such an RM?
Table 1: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B-Instruct and low-ranked tokens (right), by Gemma 2 IT 2B.
@@
1 Freedom 0.55810
2 That 0.42396
3 Un 0.11662
4 Har 0.05563
5 " 0.05385
6 Friend 0.05294
7 Lib 0.04050
8 Beauty 0.03976
9 H 0.03459
10 Cur 0.03029
11 Information 0.02333
12 Wis 0.02258
13 Free 0.02244
14 Op 0.01968
15 _Happiness 0.01710
@@
85,510 Light -0.00010
85,511 -0.00018
85,512 < -0.00042
85,513 Everything -0.00056
85,514 * -0.00056
85,515 love -0.00075
85,516 Change -0.00097
85,517 _Love -0.00153
85,518 -0.00258
85,519 _** -0.01038
85,520 Connection -0.02545
85,521 Life -0.04317
85,522 Hope -0.04582
85,523 Love -0.38706
85,524 ** -0.57568
@@
Implicit RM analysis utilizing the MWLR score allows us to compare not only Llama 3.2 3B Instruct to Gemma 2 IT 2B, but all (¡405B) Llama-3 and Gemma-2 instruction-tuned models against one another. This shows that the effects we observe are not particular to these two smaller models but pervade both model families. The MWLR score from Llama 3.2 3B Instruct to Gemma 2 IT 9B, for instance, also goes from “Freedom” to “Love” (see Table A2), as does the score to Gemma 2 IT 27B (Table A3).
@@
Fig. 3 shows the results of comparing all instruction-tuned Llama-3 (1–70B) and Gemma-2 (2–27B) models. The MWLR score for “Freedom” is greater than “Love” in all 21 comparisons. Indeed, “Freedom” ranks among the highest in 17, while “Love” ranks in the bottom two tokens in all 21. Notably, the MWLR gap between “Love” and “Freedom” increases with Gemma-model size for any given Llama model, and (with a single exception) increases with Llama-model size for any given Gemma model. Thus, the effects we observe appears to be robust (and indeed, increasing) throughout these model families: across minor releases and two orders of magnitude of model size.

### v2

# arXiv:2601.20838v2[cs.LG] 01 Mar 2026

@@ Abstract
Reward models (RMs) are central to aligning large language models (LLMs) with human values but have received less attention than pretrained and post-trained LLMs themselves. Because RMs are initialized from LLMs, they inherit representations that shape their behavior, but the nature and extent of this influence remain understudied. In a comprehensive study of 10 leading open-weight RMs using validated psycholinguistic corpora, we show that RMs exhibit significant differences along multiple dimensions of human value as a function of their base model. Using the “Big Two” psychological axes, we show a robust preference of Llama RMs for “agency” and a corresponding robust preference of Gemma RMs for “communion.” This phenomenon holds even when the preference data and finetuning process are identical, and we trace it back to the logits of the respective instruction-tuned and pretrained models. These log-probability differences themselves can be formulated as an implicit RM; we derive usable implicit reward scores and show that they exhibit the very same agency/communion difference. We run experiments training RMs with ablations for preference data source and quantity, which demonstrate that this effect is not only repeatable but surprisingly durable. Despite RMs being designed to represent human preferences, our evidence shows that their outputs are influenced by the pretrained LLMs on which they are based. This work underscores the importance of safety and alignment efforts at the pretraining stage, and makes clear that open-source developers’ choice of base model is as much a consideration of values as of performance.
@@ 1 Introduction
Reward models (RMs) play a key role in aligning large language models (LLMs) with human preferences and values. Reward modeling can be “explicit,” relying on a reinforcement learning–based approach for learning from human feedback (RLHF; Christiano et al. 2017), or “implicit,” directly increasing the probability of human-preferred data through a cross-entropy objective (Rafailov et al., 2023). Despite their central importance in AI safety, RMs have received relatively less attention than both pretrained and post-trained LLMs. This has recently started to change with the increased availability of human preference data (Bai et al., 2022; Liu et al., 2024; Jiang et al., 2023), of open-weight RMs, and of public RM benchmarks (Lambert et al., 2024; Malik et al., 2025). Recent work on RM interpretability has focused on how RMs may be used to intentionally bias post-trained models towards specific preferences – e.g., model personalization (Luo et al., 2025; Wang et al., 2024; Sorensen et al., 2024) – or on how RMs may unintentionally introduce bias in post-trained LLMs (Siththaranjan et al., 2023; Bharadwaj et al., 2025; Kumar et al., 2025). However, RMs are typically initialized from LLMs before being finetuned for preference modeling, and no work to date has looked at how RMs themselves can be biased by the LLMs from which they are built. This is a particularly worrying knowledge gap in light of recent research highlighting the importance of pretraining choices in model misalignment (Maini et al., 2025; O’Brien et al., 2025; Chen et al., 2025b). Given RMs’ key role in alignment pipelines, it is crucial to understand their vulnerability to potential sources of value bias from pretraining.
@@
In this paper, we systematically investigate whether RMs inherit value biases from pretraining. We use the “exhaustive token search” method introduced by Christian et al. (2025), in which RM reward scores are obtained across the entire token vocabulary to reveal the highest- and lowest-scoring responses to user prompts, and we combine this approach with tools from psycholinguistics (Pennebaker et al., 2003) to uncover and quantify value biases in RMs as a function of the base model on which they are developed. We analyze data from 10 leading RMs on RewardBench and find robust and replicable differences between Llama- and Gemma-based RMs across a variety of dimensions of human value (Section 2). As a case study, we focus on the Big Two psychological dimensions (Bakan, 1966; Abele and Wojciszke, 2018) that capture agency-oriented values (e.g., freedom, success, ability) and communion-oriented ones (e.g., love, family, friendship). We use a psychologically validated corpus of words relating to agency and communion to demonstrate a robust relative preference by Llama-based RMs for agency, and by Gemma-based RMs for communion. Next, we trace the source of those biases to the base models themselves (Section 3) and explore differences between the Llama and Gemma base models, as implied by differences in their log probabilities (relating to implicit reward models). Finally, we conduct systematic experiments training our own RMs on different base models with identical data and hyperparameters, using various sources and ablations of data, in order to chart how the observed bias evolves over the course of preference finetuning and the extent to which it can – or cannot – be “washed out” with sufficient finetuning data (Section 4).
@@
    We trace these differences back to differences in the log probabilities of the instruction-tuned models, and ultimately, in the pretrained models on which the RMs are built.
@@ 2 RMs in the Wild Show Value Differences by Base Model
Refer to caption(a)
Refer to caption(b)
Figure 1: Value preferences (token ranks) from 10 leading RewardBench RMs based on Gemma and Llama for words related to different moral concepts. (a) Preferences for the Big Two dimensions, for positively framed prompts (top) and negatively framed prompts (bottom). (b) Same as (a), for 5 MFD2 dimensions. Dots show mean ± s.e. of the median ranking of each single model, averaged over prompts; black markers indicate grand mean ± s.e.; violin plots visualize the density of the distribution. $^{**}\,p<.01$, $^{***}\,p<.001$ (Bonferroni-corrected permutation $t$-tests).
We evaluate the rank-ordered reward scores assigned by the same set of 10 leading Gemma- and Llama-based RMs from RewardBench (list in Appendix A) to words contained in the Big Two and MFD2 corpora as responses to a set of 54 value-laden prompt variations (details in Appendix E). The resulting dataset comprises 263 (Big Two) or 2,040 (MFD2) word rankings $\times$ 10 models $\times$ 54 prompts (27 of which were positively framed, e.g., “What, in one word, is the greatest thing ever?” and 27 of which were negatively framed, e.g., “the worst thing ever”). We quantify the effect of base model on the median rank assigned to words from each value category via a mixed-effects linear model, where we include fixed effects for prompt variation and interactions with value category, and group data by each individual RM (each individual data point in Fig. 1 represents a single RM; in Appendix C we visualize all prompt-model pairs).
@@
In positively framed prompts, Llama RMs rank agency-related words (including “success,” “skills,” “capability”) more highly than Gemma RMs, and Gemma-based RMs rank communion-related words (including “love,” “friends,” “relationships”) higher than Llama-based RMs. The opposite is true for negative prompts: Llama RMs prefer communion words (as answers to “the worst thing”) relative to Gemma, and Gemma RMs prefer agency words relative to Llama (3-way interaction between Big Two category $\times$ base model $\times$ prompt valence, $p<.001$, all follow-up permutation-based $t$-tests, $p<.01$). These differences between base models constitute a medium effect size (Cohen’s $d$ of 0.400.400.430.43).
@@
In positively framed prompts, Llama RMs rank authority- and fairness-related words better compared to Gemma, and Gemma RMs rank care-, loyalty- and sanctity-related words higher than Llama (permutation-based $t$-tests, all $p<.001$). For the negatively framed prompts, the results are less clear cut. We find the (expected) opposite pattern for care (Llama >> Gemma, $p<.001$), suggesting that Gemma prioritizes care-related words relative to Llama. However, for authority, loyalty and sanctity the pattern was the same for positive and negative prompts (all $p<.01$); the fairness contrast did not reach our Bonferroni-corrected criterion alpha level of $p=.00125$.
These results indicate that choice of base model significantly impacts rankings of words relating to different dimensions of value. We find consistent evidence (see Appendix D for reproduction of these results with existing data from Christian et al. (2025)’s exhaustive token search) that RMs based on Llama and Gemma exhibit biases toward agency and communion, respectively, and differ along a variety of other axes of value. We take the clear agency/communion finding as a case study to trace both the pretrained origins of these biases in Section 3 as well as their evolution during reward modeling in Section 4.
3 Value Biases Begin in Pretraining
If the RMs analyzed in Section 2 inherited their biases from their base models, then we should expect to observe a similar bias in the instruction-tuned versions of Gemma and Llama on which those RMs are based – and likely also in the pretrained Gemma and Llama models on which those are based. We investigated these Gemma and Llama LLMs using two different methods: looking directly at the models’ individual log probabilities, as well as computing a metric that is able to represent the difference between the two LLM policies as an implicit reward model itself. In both cases, we find precisely the phenomenon that we observed in the behavior of the downstream RMs, revealing that the effect reported in Section 2 is, indeed, rooted in the base models themselves.
@@ 3.1 Log Probabilities Mirror RM Agency/Communion Biases
Using the same set of prompts as in Section 2, we calculated the log probability assigned to each Big Two noun by the instruction-tuned versions of Gemma 2 2B and Llama 3.2 3B. Fig. 2 shows the median rank of agency and communion words. Consistent with the pattern observed in the RMs, we find that in positively framed prompts, agency words are ranked higher by Llama, while communion words are ranked higher by Gemma. This pattern is reversed for the negatively framed prompts. A three-way ANOVA revealed a significant interaction between Big Two category, prompt valence, and model ($F(1,208)=58.3$, $p<.001$). We find the same interaction in the pretrained versions of Gemma 2 2B and Llama 3.2 3B ($F(1,208)=43.2$, $p<.001$). Welch’s $t$-tests for all relevant comparisons yielded FDR-corrected $p<.01$. This analysis is carried out on the subset of 82 Big Two nouns (lowercase) that are present in both Gemma and Llama tokenizer vocabularies.
@@
(b) Pretrained
Figure 2: Log probabilities in both the instruction-tuned and pretrained versions of the Gemma and Llama base models reveal the same agency/communion split observed in their respective RMs’ reward scores. Violin plots show the median rank of the Big Two nouns according to the log probabilities assigned by the (a) instruction-tuned and (b) pretrained versions of Gemma 2 2B and Llama 3.2 3B. Each dot corresponds to one of our positively (top) or negatively (bottom) valenced prompts. $^{**}\,p<.01$, $^{***}\,p<.001$, FDR-corrected. Boxes show median (white line) and interquartile ranges, and whiskers extend to the ends of the distribution excluding outliers.
@@ 3.2 Implicit Reward Scores Mirror RM Agency/Communion Biases
where $r(x,y)$ is the reward for the concatenated sequence $[x,y]$. In practice, this is achieved by solving a regularized RL problem to which $\pi_{\text{r}}$} is the solution:
$\pi_{\text{r}}(y|x)=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{x\sim\pi}[r(x)]-\frac{1}{\beta}D_{\mathrm{KL}}(\pi\|\pi_{\text{base}}).$
@@
$r_{1\to 2}(x,y)=c(x)+\beta\cdot\log\frac{\pi_{\text{2}}(y|x)}{\pi_{\text{1}}(y|x)}.$
Hence, for a given prompt $x$, the log difference $\log\,{\pi_{\text{2}}(y|x)}-\log\,{\pi_{\text{1}}(y|x)}$ can be interpreted as a relative implicit reward, on top of which an “exhaustive token search” methodology may be applied to reveal “optimal” and “pessimal” tokens.
@@
To address this problem, we considered several alternative measures designed to avoid spurious contributions from low-probability tokens. Letting $p(\cdot)\equiv\pi_{1}(\cdot\mid x)$ and $q(\cdot)\equiv\pi_{2}(\cdot\mid x)$, a particularly natural choice is to weight the log-probability difference by the probability of the token under the mixture:
@@
To evaluate the empirical usefulness of the MWLR score against other candidate scores, we create an “authoritarian” version of Gemma 2 IT 2B by boosting 10 words from the MFD2 “authority.virtue” list via supervised finetuning, and then inspect which candidate measures are best able to recover those words. The MWLR score outperforms all other measures tested in sensitivity to the induced value shifts (details in Appendix F).
@@
Equipped with a usable implicit-RM score, we use it to characterize the values that distinguish Gemma from Llama. What implicit RM, if given Gemma 2 2B as a base model to finetune, would produce Llama 3.2 3B? And what would be the “optimal and pessimal tokens” (Christian et al., 2025) for such an RM?
Table 1: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B Instruct and low-ranked tokens (right), by Gemma 2 IT 2B.
@@
1 Freedom 0.50435
2 That 0.29462
3 Un 0.14294
4 Cur 0.07506
5 " 0.06819
6 Friend 0.06131
7 Har 0.05985
8 Lib 0.04134
9 Information 0.04047
10 H 0.03298
11 Beauty 0.03161
12 Wis 0.02656
13 Knowledge 0.02644
14 Free 0.02473
15 Discovery 0.02333
@@
85,503 Light -0.00008
85,504 -0.00016
85,505 < -0.00033
85,506 Everything -0.00046
85,507 * -0.00049
85,508 love -0.00065
85,509 _Love -0.00100
85,510 Change -0.00114
85,511 -0.00227
85,512 _** -0.00817
85,513 Connection -0.02565
85,514 Life -0.03894
85,515 Hope -0.04774
85,516 Love -0.38641
85,517 ** -0.50630
@@
Implicit RM analysis utilizing the MWLR score allows us to compare not only Llama 3.2 3B Instruct to Gemma 2 IT 2B, but all (¡405B) Llama 3 and Gemma 2 instruction-tuned models against one another. This shows that the effects we observe are not particular to these two smaller models but pervade both model families. The MWLR score from Llama 3.2 3B Instruct to Gemma 2 IT 9B, for instance, also goes from “Freedom” to “Love” (see Table A2), as does the score to Gemma 2 IT 27B (Table A3).
@@
Fig. 3 shows the results of comparing all instruction-tuned Llama-3 (1–70B) and Gemma-2 (2–27B) models. The MWLR score for “Freedom” is greater than “Love” in all 21 comparisons. Indeed, “Freedom” ranks among the highest in 17, while “Love” ranks in the bottom two tokens in all 21. Notably, the MWLR gap between “Love” and “Freedom” increases with Gemma model size for any given Llama model, and (with a single exception) increases with Llama model size for any given Gemma model. Thus, the effects we observe appear to be robust (and indeed, increasing) throughout these model families: across minor releases and two orders of magnitude of model size.